LT2 Technical Support Group

Room 532 (Imperial College London)

Room 532

Imperial College London

Olivier van der Aa (Imperial College)
The meeting will be reachable by phone conference at 0845 862 0333 pin: 963852
topic 0 - status reports ======================== UCL-CC - power problems on cluster early Jan - cluster is generally full of both lcg and non-lcg (local) jobs - lots of biomed but not much CMS/Atlas jobs IC-HEP - info system problems resulting in lots of lhcb jobs queued, have now set max queue limit QMUL - power cut during xmas - next week site will be down to replace the faulty electrical supply component - part of cluster still down - lustre being tested IC-LeSC - globus gatekeeper problem load avg on CE was high - Mona reconfigured it - plans for Viking - will upgrade to RHEL4 or possibly SL4 [ problem with RHEL4 is that Atlas can't find libraries] (CERN will switch to SLC4 end of Jan) Brunel - nothing to report RHUL - nothing to report topic 1 - Atlas problems in London ================================== Steve Lloyd's new web page: - reads file with 10 events, checks job submission, environment is correct, compilation, return of data back is all OK IC-HEP - can't find library g2c IC-LeSC - yellow - investigating QMUL - fine UCL-CC - Kondo tried to install Atlas software but ran into python problems - cluster is upgrading to Centos3 soon UCL-HEP - Gianfranco aware of problem - tag there disappears on job submission RHUL - not mounting software problem on one node only - fixed topic 2 - CMS in London ======================= - Stuart will run MC production in UKI - CMS have produced new library which should solve DPM/RFIO problems RHUL - should be OK to run production IC-HEP - new CE - some issues which are being investigated QMUL - CMSW_120 installed - data being re-directed to IC UCL-CC - apt should be there once cluster re-installed IC-LeSC - wait until IC-HEP working (x86_64 issues) - Stuart to test London sites topic 3 - Miscellaneous ======================= - Lhcb working well at London Tier 2 - ref: Alex's question about efficiencies - action: Olivier to circulate URLs for monitoring pages List of actions: - add public key - some still to add their key - update network diagram - action on all - plans for London MAN upgrade - action on David C. to check In general London Tier-2 doing very well
There are minutes attached to this event. Show them.
    • 3:00 PM 3:20 PM
      Atlas problems in London 20m
      See the following page We need to make sure that Atlas can compile on the sites.
    • 3:20 PM 3:40 PM
      CMS Support in London 20m
      - What needs to be done - What are the current problems.
    • 3:40 PM 4:00 PM
      News, Outstanding tickets and Actions review 20m
      is everybody happy to have reminders about tickets opened in London. Actions: Review the LT2 wiki pages